# II. Carbon-enhanced metal-poor starsthanks: All M3DIS 3D RHD CEMP models are available under https://nlte.mpia.de

### v1

@@ 3.4 Galactic CEMP fraction in 3D
To quantify the consequences for Galactic fractions of CEMP-r/s and CEMP-no sub-groups, splitting the CEMP sample based on the neutron capture abundances is necessary. However, for a significant portion of the SAGA sample no Ba or Eu measurements are available. This issue has been discussed by Yoon et al. (2016), who derived an alternative classification based on $\rm A(C)$ alone, for which they claim a success rate of $\rm 87\%$ for CEMP-r/s classification and $\rm 93\%$ for CEMP-no stars. In their method, stars are associated with the CEMP-r/s group, if a high absolute carbon abundance $\rm A(C)>7.1$ is measured, while CEMP-no stars consequently have $\rm A(C)\leq 7.1$. The discussion of their approach goes beyond the boundaries of this paper. We emphasize that their CEMP subgroup association is empirical, which means the systematic offsets presented here will either lead to a re-classification of CEMP stars in the same system, if the separation of Yoon et al. (2016) is still valid, or require a change of the system through new $\rm A(C)$ limits based on new C abundances.
@@
In 1D, there is a significant fraction of over 10 $\%$ of CEMP-r/s stars across all metallicities; below $\rm[Fe/H]=-4$ we find $\sim 30\%$ of all stars in our sample are CEMP-r/s stars. The fraction of CEMP-no stars has a strong metallicity dependence. Below $\rm[Fe/H]=-5$ all stars in the sample are CEMP stars, with a share of $70\%$ CEMP-no and $30\%$ CEMP-r/s. Increasing metallicity decreases the contribution from CEMP-no stars, until $\rm[Fe/H]\leq-2$, where both contribute equally to the total number of CEMP stars ($\sim 15\%$ each).
@@
We remind the reader that it must still be explored, using large self-consistent stellar samples with 3D NLTE Ba and Eu abundances, whether or not the CEMP sub-group classification is still meaningful in 3D. Because the abundance corrections themselves depend on metallicity, it is not straightforward to assume that a separation based on a gap in the 1D $\rm A(C)$ will still be present in 3D.
@@
If one considers CEMP fractions in metallicity bins rather than as cumulative sum of stars below a given metallicity, it is straightforward to estimate the mis-classification of CEMP stars around a specific metallicity. For this purpose, we include Tab. 3 in the appendix, listing the fraction of CEMP stars using the same binning as in Fig. 7. In total, we find that $\rm\sim 30\%$ of the CEMP stars between $\rm-4.5<[Fe/H]<-3.5$ are not carbon-enhanced enough to qualify as such. Between $\rm-3.5<[Fe/H]<-2.5$ and $\rm-5.5<[Fe/H]<-4.5$ the mis-classification is still $15\%$ and $12\%$, respectively.
@@
3.5 Comparison with literature
@@
We note that Placco et al. also explored the impact of 3D on C abundances in giants. Using the 3D line formation calculations for CH by Collet et al. (2007), they estimate that negative 3D corrections, in giants with $\rm log(g)<3$, of 0.3-0.3 for $\rm-2.5<[Fe/H]<-2.0$, $\rm-0.5$ for $\rm-3.0<[Fe/H]<-2.5$, and 0.7-0.7 for $\rm[Fe/H]=-3.0$ may arise. These estimates are similar to our results, however, our analysis confirms that departures from 1D HE are also prominent in the CH G-band spectra of MS stars. We note, that the effect of C-enhancement in the 3D RHD atmosphere itself was not taken into account by Collet et al. (2007) and, hence, it was not included in the 3D estimates by Placco et al. (2014). As we show in this work, 3D effects are relevant for stars in all evolutionary phases. They thus can have a more significant impact on the overall distribution of CEMP stars across all metallicities than Placco et al. anticipated.
@@
Figure 10: Cumulative CEMP-no fraction predictions from Hartwig et al. (2018), compared to our data. Below $\rm[Fe/H]<-3.75$, the fraction of all CEMP stars in our sample are shown (circles). See text. 3D CEMP data is shown in blue (right panel), while 1D data is shown in red (left panel). The red line in the left panel corresponds to the fiducial model of Hartwig et al. (2018), which includes $40\%$ faint (from Ishigaki et al. 2014) and $60\%$ normal CCSNe (from Nomoto et al. 2013). The blue curve in the right panel corresponds to their $20\%$ faint, $80\%$ normal CCSNe model.
@@
Since we find lower CEMP fractions at all metallicities, a question arises: what are the implications of lower CEMP occurrence rates in the early Galaxy? Hartwig et al. (2018) investigated the formation of the first generation of stars using a semi-analytical model [^3] An early version of the A-SLOTH model; see Hartwig et al. (2022, 2024). based on dark matter halo merger trees that yield MW-like halos. They probed the influence of various input parameters to their simulations on the observed CEMP occurrence rate as a function of [Fe/H] and compared their predictions to the data of Placco et al. (2014). Faint SNe were considered as the primary channel for the formation of CEMP-no stars (e.g. Klessen & Glover 2023, references therein). These objects, although spanning the same mass range as normal core-collapse SNe (CCSNe), produce little 56Ni because of low explosion energy and/or effects of geometry in the explosion (Tominaga et al. 2007). Faint Pop III SNe generally produce a higher $\rm[C/Fe]$ ratio, because Fe-peak elements fall back onto the remnant and are not ejected. In the fiducial model, which is tuned to reproduce the 1D CEMP data from Placco et al. (2014), Hartwig et al. (2018) include $40\%$ of faint (Ishigaki et al. 2014), and $60\%$ normal CCSNe (Nomoto et al. 2013). In addition, Hartwig et al. (2018) provide predicted CEMP fractions for a model that includes significantly fewer faint SNe (only $20\%$ instead of $40\%$). A comparison of specifically those two models with the data presented in this work can be found in Fig. 10. We note that Hartwig et al. do not account for r-process and s-process enrichment in their model and hence only provide predictions for the fraction of CEMP-no stars. As mentioned in Sec. 3.4, the CEMP-r/s or CEMP-no classification without s- and r-process abundance measurements is a relatively uncertain endeavour. At low metallicities, the rigid A(C) cut may especially lead to a bias by wrongfully assigning high [C/Fe] group III stars as CEMP-r/s (see Fig. 8), even though group III predominantly consists of CEMP-no stars (e.g. Yoon et al. 2016). Motivated by Fig. 8, we choose to include all CEMP stars below $\rm[Fe/H]<-3.75$ – corresponding to the low-metallicity end of the blue group I ellipse – and only CEMP-no stars at higher metallicities in the comparison with Hartwig et al. (2018) shown in Fig. 10. This effectively assumes that group III is composed exclusively of CEMP-no stars. We note that this assumption does not affect the overall conclusions, because we are mainly focusing on the shift between 3D CEMP and 1D and the metallicity region above 3.5-3.5. For a more detailed investigation, s- and r-process abundance measurements are essential.
@@
Remarkably, their model with a $20\%$ faint SNe fraction (blue, solid line) is in excellent agreement with our 3D-corrected CEMP fractions (blue crosses + circles), while their fiducial model is in good agreement with our 1D data (in red). We additionally note that the systematic offset between 3D CEMP and 1D data at metallicities above 3-3 resembles the systematic offset between the fiducial and $20\%$ faint SNe Hartwig et al. models very well. At intermediate to low metallicities the CEMP-r/s – CEMP-no classification becomes less precise due to the mixing of different CEMP groups. Nevertheless, because of the higher CEMP-no contribution in 3D, the agreement between model and observations is excellent also below $\rm[Fe/H]=-3.75$, regardless of whether the assumption of nearly $100\%$ CEMP-no stars in group III holds. In 1D, this is not the case; only after including all CEMP group III stars is the agreement restored, indicating that the CEMP-no classification is the culprit here. However, independent of the agreement at low metallicity in 1D, Fig. 10 shows clearly that 3D-correction induced shifts in the CEMP fraction can be met with a decrease in the contribution of faint SNe in GCE. Hartwig et al. (2018) furthermore note that the $20\%$ faint SNe model produces the approximately same metallicity distribution function for the halo as their fiducial model, and hence remains in good agreement with observations. An overestimation of the CEMP fraction thus leads to an overestimation of the contribution from faint SNe in the evolution of the early Galaxy. This is very interesting, as faint SNe will also impact the chemical enrichment of other chemical elements and hence large robust 3D NLTE abundance samples from next-generation facilities, like 4MOST (Bensby et al. 2019), will be in position to constrain the contribution of faint SNe to the GCE.
@@ 4 Conclusions
In this work we show that physically-realistic 3D RHD models of stellar atmospheres are crucial for the determination of accurate abundances of carbon in stars. We also show that these have a direct impact on understanding nucleosynthesis and Galactic chemical evolution (GCE).
@@
We explore the effects of 3D convection on the critical diagnostic G-band in spectra of MS and sub-giant stars. The G-band arises due to absorption by the CH molecule and it provides the strongest observational constraint on C abundance and [C/Fe] ratio in metal-poor stars. To compute 3D radiation-hydrodynamics simulations of stellar sub-surface convection, we use our new code M3DIS presented in Eitner et al. (2024). The uppermost parts of these models are used to calculate detailed 3D spectrum synthesis around the G-band and to derive a grid of 3D abundance corrections to [C/Fe] for the entire range of metallicities, $6\leq$ [Fe/H] $\leq-2$, relevant for studies of GCE. The 3D models are computed for the first time using C-enhancement in line with low-metallicity observations. We compare the 3D synthetic spectra in the G-band with the equivalent simulations computed using classical 1D HE models and quantify the amplitude of 3D effects on stellar [C/Fe] abundances.
@@
We find that 3D effects on the G-band are substantial and lead to lower estimates of the carbon abundance. This effect arises due to cooler outer temperature structures of 3D RHD models, compared to 1D HE structures. Formation of molecules is much more efficient in cooler 3D models, which implies that the spectral lines of CH are stronger. As a consequence, a lower abundance of C is required to fit a given observed spectrum, and the 3D corrections are negative. Interestingly, increasing C abundance in 3D models at the same metallicity (set by iron) yields the opposite effect: the G-band weakens. Physically, this is because increasing C (to levels observed in CEMP stars) implies higher radiative opacity in the models and leads to slightly hotter 3D models, compared to their scaled-solar counterparts. The 3D CEMP A(C) corrections for MS stars may exceed $\rm A(C)_{3D\ CEMP}-A(C)_{1D}=-0.6\ dex$ at $\rm[Fe/H]=-5$. For sub-giants, the 3D CEMP effects are on the order of $\rm\sim-0.7$ dex, in good agreement with previous estimates in the literature (Collet et al. 2018; Gallagher et al. 2016, 2017). The differences between 3D and 1D A(C) decrease with increasing [Fe/H].
@@
We apply 3D CEMP A(C) abundance corrections to a sample of metal-poor stars from the SAGA database to investigate the impact on the Galactic CEMP population. We find a significant fraction of CEMP stars across all metallicities that no longer classify as C-enhanced. In the range between $\rm-4.5<[Fe/H]<-3.5$ we find $\sim 30\%$ fewer CEMP stars after applying the 3D abundance corrections, between $\rm-3.5<[Fe/H]<-2.5$ the number is reduced by $15\%$. We furthermore find that the cumulative CEMP fraction below $\rm[Fe/H]=-3$ is $\rm\sim 20\%$ lower in 3D than it appears in 1D. At lower metallicities the distribution remains mostly unchanged, because the same $\rm A(C)$ corresponds to a higher $\rm[C/Fe]$. As a consequence the $\rm A(C)$ plateau remains present in 3D, though it appears to be slightly tilted by $10^{\circ}$. When considering the CEMP-no and CEMP-r/s sub-groups based on $\rm A(C)$ (Yoon et al. 2016), we find a reduction of the CEMP-r/s sample size in 3D, as a significant number of stars are either shifted below the CEMP limit of $\rm[C/Fe]=0.7$, or below the CEMP-r/s limit of $\rm A(C)=7.1$. Consequently, the use of 3D-corrected C abundances leads to a higher fraction of CEMP-no stars, as some stars previously classified as CEMP-r/s are re-classified. Despite their generally lower metallicities, CEMP-no stars retain a high $\rm A(C)$, ensuring that most remain classified as carbon-enhanced.
@@
The fraction of CEMP stars as a function of metallicity is a tracer of the contribution of faint SNe to the GCE (Hartwig et al. 2018; Klessen & Glover 2023). Following the models of Hartwig et al. (2018), we find that a lower fraction (only 20%) of faint SNe in the Pop III galaxy formation models can describe the 3D CEMP distributions. This is an interesting result that remains to be confirmed with detailed GCE models and other chemical elements, which are sensitive to faint SNe contribution.

### v2

@@ 3.4 Galactic CEMP fraction in 3D
To quantify the consequences for Galactic fractions of CEMP-r/s and CEMP-no subgroups, splitting the CEMP sample based on the neutron capture abundances is necessary. However, for a significant portion of the SAGA sample no Ba or Eu measurements are available. This issue has been discussed by Yoon et al. (2016), who derived an alternative classification based on $\rm A(C)$ alone, for which they claim a success rate of $\rm 87\%$ for CEMP-r/s classification and $\rm 93\%$ for CEMP-no stars. In their method, stars are associated with the CEMP-r/s group, if a high absolute carbon abundance $\rm A(C)>7.1$ is measured, while CEMP-no stars consequently have $\rm A(C)\leq 7.1$. The discussion of their approach goes beyond the boundaries of this paper. We emphasise that their CEMP subgroup association is empirical, which means the systematic offsets presented here will either lead to a re-classification of CEMP stars in the same system, if the separation of Yoon et al. (2016) is still valid, or require a change of the system through new $\rm A(C)$ limits based on new C abundances.
@@
In 1D, there is a significant fraction of over 10 $\%$ of CEMP-r/s stars across all metallicities; below $\rm[Fe/H]=-4$ we find $\sim 30\%$ of all stars in our sample are CEMP-r/s stars. The fraction of CEMP-no stars has a strong metallicity dependence. Below $\rm[Fe/H]=-5$ all stars in the sample are CEMP stars, with a share of $70\%$ CEMP-no and $30\%$ CEMP-r/s. Increasing metallicity decreases the contribution from CEMP-no stars, until $\rm[Fe/H]\leq-2$, where the two contribute equally to the total number of CEMP stars ($\sim 15\%$ each).
@@
We remind the reader that it must still be explored, using large self-consistent stellar samples with 3D NLTE Ba and Eu abundances, whether or not the CEMP subgroup classification is still meaningful in 3D. Because the abundance corrections themselves depend on metallicity, it is not straightforward to assume that a separation based on a gap in the 1D $\rm A(C)$ will still be present in 3D.
@@
If one considers CEMP fractions in metallicity bins rather than as cumulative sum of stars below a given metallicity, it is straightforward to estimate the misclassification of CEMP stars around a specific metallicity. Table 3 lists the fraction of CEMP stars using the same binning as in Fig. 7. In total, we find that $\rm\sim 30\%$ of the CEMP stars in the range $\rm-4.5<[Fe/H]<-3.5$ are not carbon-enhanced enough to qualify as such. In the ranges $\rm-3.5<[Fe/H]<-2.5$ and $\rm-5.5<[Fe/H]<-4.5,$ the misclassification is still $15\%$ and $12\%$, respectively.
@@
3.5 Comparison with the literature
@@
We note that Placco et al. also explored the impact of 3D on C abundances in giants. Using the 3D line formation calculations for CH by Collet et al. (2007), they estimate that negative 3D corrections, in giants with $\rm log(g)<3$, of 0.3-0.3 for $\rm-2.5<[Fe/H]<-2.0$, $\rm-0.5$ for $\rm-3.0<[Fe/H]<-2.5$, and 0.7-0.7 for $\rm[Fe/H]=-3.0,$ may arise. These estimates are similar to our results; however, our analysis confirms that departures from 1D HE are also prominent in the CH G-band spectra of MS stars. We note that the effect of C-enhancement in the 3D RHD atmosphere itself was not taken into account by Collet et al. (2007) and, hence, was not included in the 3D estimates by Placco et al. (2014). As we show in this work, 3D effects are relevant for stars in all evolutionary phases. They thus can have a more significant impact on the overall distribution of CEMP stars across all metallicities than Placco et al. anticipated.
@@
Figure 10: Cumulative CEMP-no fraction predictions from Hartwig et al. (2018) compared to our data. Below $\rm[Fe/H]<-3.75$, all CEMP stars in our sample are shown (circles). 3D CEMP data are shown in blue (right panel) and 1D data in red (left panel). The red line in the left panel corresponds to the fiducial model of Hartwig et al. (2018), which comprises $40\%$ faint (from Ishigaki et al. 2014) and $60\%$ normal CCSNe (from Nomoto et al. 2013). The blue curve in the right panel corresponds to their $20\%$ faint, $80\%$ normal CCSN model.
@@
Since we find lower CEMP fractions at all metallicities, a question arises: what are the implications of lower CEMP occurrence rates in the early Galaxy? Hartwig et al. (2018) investigated the formation of the first generation of stars using a semi-analytical model [^3] An early version of the A-SLOTH model; see Hartwig et al. (2022, 2024). based on dark matter halo merger trees that yield Milky Way-like halos. They probed the influence of various input parameters to their simulations on the observed CEMP occurrence rate as a function of [Fe/H] and compared their predictions to the data of Placco et al. (2014). Faint SNe were considered as the primary channel for the formation of CEMP-no stars (e.g. Klessen & Glover 2023, references therein). These objects, although spanning the same mass range as normal CCSNe, produce little 56Ni because of low explosion energy and/or effects of geometry in the explosion (Tominaga et al. 2007). Faint Pop III SNe generally produce a higher $\rm[C/Fe]$ ratio, because Fe-peak elements fall back onto the remnant and are not ejected. In the fiducial model, which is tuned to reproduce the 1D CEMP data from Placco et al. (2014), Hartwig et al. (2018) include $40\%$ of faint (Ishigaki et al. 2014), and $60\%$ normal CCSNe (Nomoto et al. 2013). In addition, Hartwig et al. (2018) provide predicted CEMP fractions for a model that includes significantly fewer faint SNe (only $20\%$ instead of $40\%$). A comparison of specifically those two models with the data presented in this work can be found in Fig. 10. We note that Hartwig et al. do not account for r-process and s-process enrichment in their model and hence only provide predictions for the fraction of CEMP-no stars. As mentioned in Sect. 3.4, the CEMP-r/s or CEMP-no classification without s- and r-process abundance measurements is a relatively uncertain endeavour. At low metallicities, the rigid A(C) cut especially can result in a bias by wrongfully assigning high [C/Fe] group III stars as CEMP-r/s (see Fig. 8), even though group III predominantly consists of CEMP-no stars (e.g. Yoon et al. 2016). Motivated by Fig. 8, we chose to include all CEMP stars below $\rm[Fe/H]<-3.75$ – corresponding to the low-metallicity end of the blue group I ellipse – and only CEMP-no stars at higher metallicities in the comparison with Hartwig et al. (2018) shown in Fig. 10. This effectively assumes that group III is composed exclusively of CEMP-no stars. We note that this assumption does not affect the overall conclusions, because we are mainly focusing on the shift between 3D CEMP and 1D and the metallicity region above 3.5-3.5. For a more detailed investigation, s- and r-process abundance measurements are essential.
@@
Remarkably, their model with a $20\%$ faint SN fraction (solid blue line) is in excellent agreement with our 3D-corrected CEMP fractions (blue crosses + circles), while their fiducial model is in good agreement with our 1D data (in red). We additionally note that the systematic offset between 3D CEMP and 1D data at metallicities above 3-3 resembles the systematic offset between the fiducial and $20\%$ faint SN Hartwig et al. models very well. At intermediate to low metallicities the CEMP-r/s – CEMP-no classification becomes less precise due to the mixing of different CEMP groups. Nevertheless, because of the higher CEMP-no contribution in 3D, the agreement between model and observations is excellent also below $\rm[Fe/H]=-3.75$, regardless of whether the assumption of nearly $100\%$ CEMP-no stars in group III holds. In 1D, this is not the case; only after including all CEMP group III stars is the agreement restored, indicating that the CEMP-no classification is the culprit here. However, independent of the agreement at low metallicity in 1D, Fig. 10 shows clearly that 3D-correction induced shifts in the CEMP fraction can be met with a decrease in the contribution of faint SNe in GCE. Hartwig et al. (2018) furthermore note that the $20\%$ faint SN model produces the approximately same metallicity distribution function for the halo as their fiducial model, and hence remains in good agreement with observations. An overestimation of the CEMP fraction thus leads to an overestimation of the contribution from faint SNe in the evolution of the early Galaxy. This is very interesting, as faint SNe will also impact the chemical enrichment of other chemical elements and hence large robust 3D NLTE abundance samples from next-generation facilities, like 4MOST (Bensby et al. 2019), will be in position to constrain the contribution of faint SNe to the GCE.
@@ 4 Conclusions
In this work we show that physically realistic 3D RHD models of stellar atmospheres are crucial for the determination of accurate abundances of carbon in stars. We also show that they have a direct impact on our understanding of nucleosynthesis and the GCE.
@@
We explore the effects of 3D convection on the critical diagnostic G band in spectra of MS and sub-giant stars. The G band arises due to absorption by the CH molecule and provides the strongest observational constraint on C abundances and [C/Fe] ratios in metal-poor stars. To compute 3D RHD simulations of stellar sub-surface convection, we used our new code, M3DIS, which was presented in Eitner et al. (2024). The uppermost parts of these models were used to calculate detailed 3D spectrum synthesis around the G band and to derive a grid of 3D abundance corrections to [C/Fe] for the entire range of metallicities relevant for studies of GCE, $6\leq$ [Fe/H] $\leq-2$. The 3D models were computed for the first time using C-enhancements in line with low-metallicity observations. We compared the 3D synthetic spectra in the G band with the equivalent simulations computed using classical 1D HE models and quantified the amplitude of 3D effects on stellar [C/Fe] abundances.
@@
We find that 3D effects on the G band are substantial and lead to lower carbon abundance estimates. This effect arises due to cooler outer structures of 3D RHD models compared to 1D HE structures. The formation of molecules is much more efficient in cooler 3D models, which implies that the spectral lines of CH are stronger. As a consequence, a lower abundance of C is required to fit a given observed spectrum, and the 3D corrections are negative. Interestingly, increasing C abundances in 3D models at the same metallicity (set by iron) yields the opposite effect: the G band weakens. Physically, this is because increasing C (to levels observed in CEMP stars) implies a higher radiative opacity in the models and leads to slightly hotter 3D models compared to their scaled-solar counterparts. The 3D CEMP A(C) corrections for MS stars can exceed $\rm A(C)_{3D\ CEMP}-A(C)_{1D}=-0.6\ dex$ at $\rm[Fe/H]=-5$. For sub-giants, the 3D CEMP effects are of the order of $\rm\sim-0.7$ dex, in good agreement with previous estimates in the literature (Collet et al. 2018; Gallagher et al. 2016, 2017). The differences between 3D and 1D A(C) decrease with increasing [Fe/H].
@@
We applied 3D CEMP A(C) abundance corrections to a sample of metal-poor stars from the SAGA database to investigate the impact on the Galactic CEMP population. We find a significant fraction of CEMP stars across all metallicities that are no longer classified as C-enhanced. In the range $\rm-4.5<[Fe/H]<-3.5$ we find $\sim 30\%$ fewer CEMP stars after applying the 3D abundance corrections, and in the range $\rm-3.5<[Fe/H]<-2.5$ the number is reduced by $15\%$. We furthermore find that the cumulative CEMP fraction below $\rm[Fe/H]=-3$ is $\rm\sim 20\%$ lower in 3D than it appears in 1D. At lower metallicities, the distribution remains mostly unchanged because the same $\rm A(C)$ corresponds to a higher $\rm[C/Fe]$. As a consequence, the $\rm A(C)$ plateau remains present in 3D, though it appears to be tilted by $10^{\circ}$. When considering the CEMP-no and CEMP-r/s subgroups based on $\rm A(C)$ (Yoon et al. 2016), we find a reduction in the CEMP-r/s sample size in 3D, as a significant number of stars are shifted either below the CEMP limit of $\rm[C/Fe]=0.7$ or below the CEMP-r/s limit of $\rm A(C)=7.1$. Consequently, the use of 3D-corrected C abundances leads to a higher fraction of CEMP-no stars, as some stars previously classified as CEMP-r/s are re-classified. Despite their generally lower metallicities, CEMP-no stars retain a high $\rm A(C)$, ensuring that most remain classified as carbon-enhanced.
@@
The fraction of CEMP stars as a function of metallicity is a tracer of the contribution of faint SNe to the GCE (Hartwig et al. 2018; Klessen & Glover 2023). Following the models of Hartwig et al. (2018), we find that a lower fraction (only 20%) of faint SNe in the Pop III galaxy formation models can describe the 3D CEMP distributions. This is an interesting result that remains to be confirmed with detailed GCE models and other chemical elements, which are sensitive to faint SN contributions.